# Vision Transformers Don’t Need Trained Registers

### v4

@@ Abstract
We investigate the mechanism underlying a previously identified phenomenon in Vision Transformers – the emergence of high-norm tokens that lead to noisy attention maps (Darcet et al., 2024). We observe that in multiple models (e.g., CLIP, DINOv2), a sparse set of neurons is responsible for concentrating high-norm activations on outlier tokens, leading to irregular attention patterns and degrading downstream visual processing. While the existing solution for removing these outliers involves retraining models from scratch with additional learned register tokens, we use our findings to create a training-free approach to mitigate these artifacts. By shifting the high-norm activations from our discovered register neurons into an additional untrained token, we can mimic the effect of register tokens on a model already trained without registers. We demonstrate that our method produces cleaner attention and feature maps, enhances performance over base models across multiple downstream visual tasks, and achieves results comparable to models explicitly trained with register tokens. We then extend test-time registers to off-the-shelf vision-language models to improve their interpretability. Our results suggest that test-time registers effectively take on the role of register tokens at test-time, offering a training-free solution for any pre-trained model released without them. [^1] Project Page: https://avdravid.github.io/test-time-registers
@@ 1 Introduction
Vision Transformers (ViTs) (Dosovitskiy et al., 2021) have become a dominant architecture in computer vision, offering strong performance across a wide range of tasks (Oquab et al., 2024; Radford et al., 2021). Recently, Darcet et al. (2024) observed a surprising property in these models: the emergence of high-norm intermediate tokens at seemingly random locations in the image during the internal computation of the ViT (see "Original" column in Figure 1). These outlier tokens were shown to appear in low-information image areas (e.g., uniform background patches) and were demonstrated to capture global image information.
@@
In this work, we argue that while registers are indeed useful, the models don’t need to be retrained with them. Instead, we show that registers can be added post hoc, without any additional training. To do this, we first investigate the mechanism underlying the emergence of high-norm tokens. We identify a sparse set of neurons – register neurons – that create the outlier tokens by contributing high-norm values to them. By directly editing the activation maps of the register neurons during the forward pass of the network, we can induce the formation of the high-norm activations at arbitrary token positions (see Figure 1). We use it to create new register tokens at test-time, even in models that were never trained with them, by appending new tokens and activating the register neurons in their positions.
We show that models with test-time registers provide comparable performance to models with trained registers on various downstream tasks (e.g., classification, segmentation, and depth prediction) and largely improve over models without registers for unsupervised object discovery (20-point correct localization improvement) and attention-based zero-shot segmentation (+5 mIOU). Next, we demonstrate that test-time registers reduce high-norm artifacts in vision-language models, improving the alignment between textual outputs and relevant visual regions. Finally, we leverage the emergent property of outlier tokens to mask local information, shifting their positions to appear on adversarial regions and improving robustness to typographic attacks by 43%.
@@
  • We determine the mechanism behind the creation of high-norm tokens in ViTs by finding register neurons that, when activated, cause the appearance of these outliers (Section 3.1).
  • We demonstrate that activating the register neurons at test-time in other image locations shifts the high-norms to the corresponding tokens (Section 3.2).
  • We present a training-free method for adding registers to models that were trained without them, by appending additional tokens and activating register neurons in their positions (Section 4).
  • We evaluate the performance of models with test-time registers and show that it is comparable to models with trained registers, thus eliminating the need for retraining models with registers from scratch (Section 5).
Figure 1: Controlling high-norm tokens in Vision Transformers. As shown in Darcet et al. (2024), high-norm outlier tokens emerge in ViTs and lead to noisy attention maps (“Original”). By identifying the mechanism responsible for their emergence, we demonstrate that we can shift these outlier tokens to arbitrary positions at test time (“Shifted”). Shifting the outlier tokens outside of the image area mimics register behavior at test-time (“w/ Test-time Register”), resulting in more interpretable attention patterns and downstream performance comparable to models retrained with registers.
@@ 2 Related work
Feature visualization in vision models. Visualizing features of computer vision models has been used for diagnostics long before the transition of the field to deep-learning (e.g., Vondrick et al. (2013)). Features in early CNN-based models were visualized to interpret their emergent computation (Zeiler & Fergus, 2014; Bau et al., 2017) and to approximate saliency maps (Itti et al., 2002). In ViTs, the attention map from the CLS token has been used for various attribution methods (Caron et al., 2021; Chefer et al., 2021), and has also been steered to improve model performance (Shi et al., 2023). Darcet et al. (2024) showed that the newer ViT-based models (Oquab et al., 2024; Radford et al., 2021) exhibit artifacts in their attentions, affecting their applicability for visualization and downstream tasks. These artifacts were connected to high-norm tokens that emerge in the transformers.
High-norm tokens in transformers. Transformers tend to create high-norm tokens during their internal computation, when trained on language tasks (Xiao et al., 2024) or on vision tasks (Darcet et al., 2024). For language models, Xiao et al. (2024) showed that transformers allocate excessive attention to these high-norm tokens, and named them “attention sinks.” Sun et al. (2024) demonstrated that “attention sinks” emerge due to previous massive neuron activations. Yona et al. (2025) linked the emergence of “attention sinks” to the inability of language models to repeatedly generate a single token, and suggested a test-time fix by zeroing out the relevant activated neurons. For vision models, Darcet et al. (2024) demonstrated the emergence of high-norm tokens in low-informative areas and suggested retraining the model with extra tokens (“registers”) to remove these high-norm tokens from the image patches. Wang et al. (2024) used a simpler fine-tuning approach in DINOv2 to avoid complete re-training. More similar to our training-free approach, Nakamura et al. (2024) suppressed artifacts during inference by modifying last-layer attention, showing that it improved image clustering performance. In contrast, we remove high-norm outliers at their source by editing the activations of the neurons that create them, aiming to eliminate their effect on later computations.
@@ 3 Interpreting the outlier tokens mechanism in ViTs
As shown by Darcet et al. (2024), high-norm outlier patches emerge during inference in various pre-trained ViTs. These patches strongly draw attention from the CLS token, resulting in artifacts in the attention maps (Figure 1). Outlier patches appear primarily in areas that exhibit high similarity to their neighboring areas (e.g., uniform background patches). Moreover, they were shown to capture global image information and lose their local patch contents (pixel + position). In this section, we study how the outlier tokens emerge in ViTs during the forward pass and discover a sparse set of neurons whose activations dictate the location of outliers. We then show that we can ablate these neurons to move the outlier positions. Our main analysis is applied to OpenCLIP ViT-B/16 (Cherti et al., 2023), with similar findings on DINOv2-L/14 (Oquab et al., 2024) shown in Section A.7.
@@ 3.1 How do outliers emerge in ViTs?
Outlier patches appear after MLPs. To identify the transformer component most responsible for outlier patches, we track the maximum norm of image patches after attention blocks and MLPs across 1000 ImageNet images (Deng et al., 2009). Figure 2 shows that outliers appear after the MLP of layer 6 in OpenCLIP. We also measure the maximum weight the CLS token attends to any patch across all heads in each layer and observe that high attention occurs after the layer 6 MLP. This observation suggests that the MLP increases the norms of certain patches, consequently creating attention sinks.
Figure 2: Outlier patches appear after MLPs; attention sinks appear after outlier patches. Left: Max norms across image patches (OpenCLIP ViT-B/16). Right: max attention scores of the CLS token in the last layer. In both plots, we average across 1000 images. The outlier norms and attention sinks occur in consecutive layers.
A small subset of neurons shows consistently high activations before outlier patches. Given that outliers appear at layer 6, we evaluate the norm contribution of individual neurons. In Figure 3, we compute the average activation of all MLP neurons in layer 6 across 1000 images, comparing the top outlier patch with a randomly selected non-outlier patch. The distribution is heavily skewed for outlier patches and more symmetric for randomly selected non-outlier patches. We find that <10 neurons consistently have large activations and appear to drive outlier formation.
Figure 3: Neuron activation distributions differ between outlier and non-outlier patches. We average the neuron activations at layer 6 for the top outlier patch and a randomly selected non-outlier patch. Non-outlier patches have a more symmetric distribution (left), whereas outlier patches show a skewed distribution with most activations near zero. A small subset of neurons (<10) consistently exhibit high activations for outlier patches across images (right).
Highly activated neurons activate on all outlier locations. We present the full activation maps of three neurons with high activations on the top outlier in Figure 4, with more shown in Section A.4. These neurons sparsely activate on all outlier positions, not just the maximum outlier position. This suggests that they are not position-specific but, rather, are responsible for outliers generally. Given these observations, we develop a method to automatically detect these neurons in the next section.
Figure 4: Highly activated neurons on the top outlier activate on all outlier positions. We present activation maps of three neurons from layer 6 that activate highly on the top outlier patch (Figure 3). These maps near-perfectly align with the high-norm outliers ("patch norms").
@@ 3.2 Register neurons
Detecting register neurons. Based on our hypothesis, we formulate a simple algorithm to find register neurons in Algorithm 1. Our method finds neurons whose average activation at outlier positions is consistently high across images. To detect outlier positions (FindOutliers), we follow Darcet et al. (2024) and output the positions for which the norms of the corresponding image tokens are above a predefined threshold. Our algorithm searches for neurons in layers before outliers start, set by the $top\_layer$ parameter. We output $top\_k$ neurons with the highest average activations. For OpenCLIP, we set $top\_layer=5$, the outlier threshold at 75, and $top\_k=10$. Since register neurons set where outliers emerge, we can also intervene upon them to move outliers to arbitrary positions.
Moving outliers with register neurons. To demonstrate the importance of these register neurons, we can use them to “move” outliers to different image locations. We apply FindRegisterNeurons to detect the register neurons and modify their activation pattern during the forward pass. More specifically, for each register neuron, we copy the highest neuron activation across the tokens into the locations of the tokens to which we want to move the outliers. We zero out the activations of the neuron elsewhere.
@@
Input: image set =$\mathcal{I}=\{I_{1},\dots,I_{M}\}$, maximum layer index $top\_layer$, number of neurons to return $top\_k$, number of neurons per layer $N$
Output: set of the top-$k$ “register neurons”
@@
2:for all $I_{i}\in\mathcal{I}$ do $\triangleright$ loop over images
3:     $O\leftarrow\textsc{FindOutliers}(I_{i})$triangleright indices of top-norm patches
4:     for all $l\in[0,...,top\_layer]$ do
5:         for all $n\in[0,...,N-1]$ do
[…]
7:         end for
8:     end for
@@
10:return $top\_k$ neurons with largest avg_act
Register neurons causally set the position of outliers. To test our ability to move outliers to arbitrary spots, we assign the highest activation to a random patch and measure its last-layer output norm, the maximum norm of other patches, and the highest last-layer CLS attention, both to the selected outlier and any other patch. Successful interventions yield high norms and attention for the targeted patch, with low values elsewhere. As shown in Figure 5, modifying the activations of register neurons effectively controls where outliers emerge. In contrast, intervening on random neurons has little effect (Section A.6). Figure 1 demonstrates that intervening on register neurons can make outliers appear in various counts and spatial patterns (e.g. a heart). We use this technique to prevent typographic attacks (Section 5.5) and for mimicking registers, as follows in the next section.
Figure 5: Intervening on activations of register neurons effectively shifts outliers to random patches and test-time registers. For all register neurons, we copy their highest activation into a selected patch and zero out the activations elsewhere. Left: norm of chosen random patch (yellow) and max norm of any other patch (blue). Right: CLS attention to chosen random patch (yellow) and max CLS attention (blue) to any other patch. Our intervention can shift outliers to randomly selected patches as well as test-time registers (see Section A.5).
@@ 4 Adding registers at test-time
Moving outliers to added tokens. As outlier patches lose their local patch information (Darcet et al., 2024), it is undesirable to have them within the image since it may affect downstream performance. Previous work has suggested retraining ViTs with extra tokens to remove high-norm artifacts and attention sinks. However, retraining existing ViTs is expensive, so we propose to add an extra input token and move the outliers there with register neurons. Importantly, these outliers are essential to the model’s internal computations and must appear either within a token or as an attention bias (Section A.11). For instance, zeroing out the register neurons’ activation maps instead of moving the outliers causes OpenCLIP ViT-B/16’s zero-shot ImageNet performance to drop $\sim$20%.
Figure 6: Qualitative results on attention maps w/ test-time registers. We produce the mean CLS attention maps of the last layer in DINOv2 and compare them to the model with trained registers. Test-time registers produce similarly high-quality maps as trained registers.
@@
Test-time register initialization. We initialize our extra token to be a vector of zeros. We assess several initialization strategies, but we find that they do not significantly affect the ability of test-time registers to store the high norms (Section A.2). Additionally, we focus our investigation on using one test-time register and report the impact of using more registers in Section A.1.
Outliers move outside the image after adding test-time registers. To evaluate whether test-time registers can absorb the outliers, we apply our intervention and measure the max norms of image patch outputs and the test-time register. Our intervention results are nearly identical to the distribution of patch norms and CLS attention after shifting outliers to random patches, previously shown in Figure 5 (see results for test-time registers in Section A.5). The image patches no longer have outliers, whereas the test-time register absorbs the outlier norms. This change is also present in the attention maps (Figure 6), which no longer have noisy artifacts and match the quality of attention maps from models with trained registers. [^2] As there is no open-source OpenCLIP trained with registers, we only present comparisons on DINOv2. See Section A.3 for attention maps in all model layers.
Test-time registers hold global information. We verify that test-time registers encapsulate global image information (e.g., image class) similarly to trained registers (Darcet et al., 2024). To assess this, we perform linear probing on both trained and test-time registers for classification on ImageNet (Deng et al., 2009), CIFAR-10, and CIFAR-100 (Krizhevsky et al., 2009). We also compare their performance to the CLS token and a token that corresponds to a central patch in the image. As shown in Table 1, the classification accuracy of test-time registers closely matches that of trained registers and is slightly lower than the CLS token performance. These results suggest that test-time registers, like their learned counterparts, effectively capture global image-level information.

### v5

@@ Abstract
We investigate the mechanism underlying a previously identified phenomenon in Vision Transformers – the emergence of high-norm tokens that lead to noisy attention maps (Darcet et al., 2024). We observe that in multiple models (e.g., CLIP, DINOv2), a sparse set of neurons is responsible for concentrating high-norm activations on outlier tokens, leading to irregular attention patterns and degrading downstream visual processing. While the existing solution for removing these outliers involves retraining models from scratch with additional learned register tokens, we use our findings to create a training-free approach to mitigate these artifacts. By shifting the high-norm activations from our discovered register neurons into an additional untrained token, we can mimic the effect of register tokens on a model already trained without registers. We demonstrate that our method produces cleaner attention and feature maps, enhances performance over base models across multiple downstream visual tasks, and achieves results comparable to models explicitly trained with register tokens. We then extend test-time registers to off-the-shelf vision-language models, yielding cleaner attention-based, text-to-image attribution. Finally, we outline a simple mathematical model that reflects the observed behavior of register neurons and high norm tokens. Our results suggest that test-time registers effectively take on the role of register tokens at test-time, offering a training-free solution for any pre-trained model released without them. [^1] Project Page: https://avdravid.github.io/test-time-registers
@@ 1 Introduction
Vision Transformers (ViTs) (Dosovitskiy et al., 2021) have become a dominant architecture in computer vision, offering strong performance across a wide range of tasks (Oquab et al., 2024; Radford et al., 2021). Recently, Darcet et al. (2024) observed a surprising property in these models: the emergence of high-norm intermediate tokens at seemingly random locations in the image during the internal computation of the ViT (see "Original" column in Figure˜1). These outlier tokens were shown to appear in low-information image areas (e.g., uniform background patches) and were demonstrated to capture global image information.
@@
In this work, we argue that while registers are indeed useful, models don’t need to be retrained with them. Instead, we show that registers can be added post hoc, without any additional training. To do this, we first investigate the mechanism underlying the emergence of high-norm tokens. We identify a sparse set of neurons – register neurons – that create the outlier tokens by contributing high-norm values to them. By directly editing the activation maps of the register neurons during the forward pass of the network, we can induce the formation of the high-norm activations at arbitrary token positions (Figure˜1). We use it to create new register tokens at test-time, even in models that were never trained with them, by appending new tokens and activating the register neurons in their positions.
We show that models with test-time registers provide comparable performance to models with trained registers on various downstream tasks (e.g., classification, segmentation, and depth prediction) and largely improve over models without registers for unsupervised object discovery (20-point correct localization improvement) and attention-based zero-shot segmentation (+5 mIOU). Next, we demonstrate that test-time registers reduce high-norm artifacts in vision-language models, improving the alignment between textual outputs and relevant visual regions. Finally, we present a simple mathematical model capturing the observed behavior of register neurons and high-norm tokens, empirically finding that test-time attention biases can mitigate outliers and attention sinks without any additional tokens.
@@
  • We determine the mechanism behind the creation of high-norm tokens in ViTs by finding register neurons that, when activated, cause the appearance of these outliers (Section˜3.1). A simple mathematical model captures the observed behavior of these neurons (Section˜6).
  • We demonstrate that activating the register neurons at test-time in other image locations shifts the high-norms to the corresponding tokens (Section˜3.2).
  • We present a training-free method for adding registers to models that were trained without them, by appending additional tokens and activating register neurons in their positions (Section˜4).
  • We evaluate the performance of models with test-time registers and show that it is comparable to models with trained registers, thus eliminating the need for retraining models with registers from scratch (Section˜5).
Figure 1: Controlling high-norm tokens in Vision Transformers. As shown in Darcet et al. (2024), high-norm outlier tokens emerge in ViTs and lead to noisy attention maps (“Original”). By identifying the mechanism responsible for their emergence, we demonstrate that we can shift these outlier tokens to arbitrary positions at test time (“Shifted”). Shifting the outlier tokens outside of the image mimics register behavior at test-time (“w/ Test-time Register”), resulting in more interpretable attention patterns and downstream performance comparable to models retrained with registers.
@@ 2 Related work
Feature visualization in vision models. Visualizing features of computer vision models has been used for diagnostics long before the transition of the field to deep-learning (e.g., Vondrick et al. (2013)). Features in early CNN-based models were visualized to interpret their emergent computation (Zeiler & Fergus, 2014; Bau et al., 2017) and to approximate saliency maps (Itti et al., 2002). In ViTs, the attention map from the [CLS] token has been used for various attribution methods (Caron et al., 2021; Chefer et al., 2021), and has also been steered to improve model performance (Shi et al., 2023). Darcet et al. (2024) showed that the newer ViT-based models (Oquab et al., 2024; Radford et al., 2021) exhibit artifacts in their attentions, affecting their applicability for visualization and downstream tasks. These artifacts were connected to high-norm tokens that emerge in the Transformers.
High-norm tokens in Transformers. Transformers tend to create high-norm tokens during their internal computation, when trained on language tasks (Xiao et al., 2024) or on vision tasks (Darcet et al., 2024). For language models, Xiao et al. (2024) showed that Transformers allocate excessive attention to these high-norm tokens, and named them “attention sinks.” Sun et al. (2024) demonstrated that “attention sinks” emerge due to previous massive activations in the residual stream. Yona et al. (2025) linked the emergence of “attention sinks” to the inability of language models to repeatedly generate a single token, and suggested a test-time fix by zeroing out the relevant activated neurons. For vision models, Darcet et al. (2024) demonstrated the emergence of high-norm tokens in low-informative areas and suggested retraining the model with extra tokens (“registers”) to remove these large norms from the image patches. Wang et al. (2024) used a simpler fine-tuning approach in DINOv2 to avoid complete re-training. Nakamura et al. (2024) suppressed artifacts during inference by modifying last-layer attention, showing that it improved image clustering performance. In contrast, we remove high-norm outliers at their source by editing the activations of the neurons that create them, aiming to eliminate their effect on later computations.
@@ 3 Interpreting the outlier tokens mechanism in ViTs
As shown by Darcet et al. (2024), high-norm outlier patches emerge during inference in various pre-trained ViTs. These patches strongly draw attention from the [CLS] token, resulting in artifacts in the attention maps (Figure˜1). Outlier patches appear primarily in areas that exhibit high similarity to their neighboring areas (e.g., uniform background patches). Moreover, they were shown to capture global image information and lose their local patch contents (pixel + position). In this section, we study how the outlier tokens emerge in ViTs during the forward pass and discover a sparse set of neurons whose activations dictate the location of outliers. We use the term "neuron" to denote a single hidden unit in the MLP layers of a Transformer block, i.e., the scalar output after the linear transformation and nonlinearity. We then show that we can edit these neurons to move the outlier positions. Our main analysis is applied to OpenCLIP ViT-B/16 (Cherti et al., 2023), with similar findings on DINOv2-L/14 (Oquab et al., 2024) shown in Section˜A.12.
@@ 3.1 How do outliers emerge in ViTs?
Outlier patches appear after MLPs. To identify the Transformer component most responsible for outlier patches, we track the maximum norm of image patches after attention blocks and MLPs across 1000 ImageNet images (Deng et al., 2009). Figure˜2 shows that outliers appear after the MLP of layer 6 in OpenCLIP. We also measure the maximum weight the [CLS] token attends to any patch across all heads in each layer and observe that high attention occurs after the layer 6 MLP. This observation suggests that the MLP increases the norms of certain patches, consequently creating attention sinks.
Figure 2: Outlier patches appear after MLPs; attention sinks appear after outlier patches. Left: Max norms across image patches (OpenCLIP ViT-B/16). Right: max attention scores of the [CLS] token in the last layer. In both plots, we average across 1000 images. The outlier norms and attention sinks occur in consecutive layers.
A small, consistent set of neurons activate highly on outlier positions. Examining layer 6, we find that just five MLP neurons exhibit consistently high contributions at the top outlier patch across images (see Section˜A.11.1). When inspecting the activation maps of three such neurons (Figure˜3, more in Section˜A.10), we observe that they sparsely activate on all outlier positions, not just the top outlier position. This suggests that these high-activating neurons are not position-specific, but rather, responsible for outliers generally. Given these observations, we develop a method to automatically detect these neurons in the next section.
Figure 3: Highly activated neurons on the top outlier activate on all outlier positions. We present activation maps of three neurons from layer 6 that activate highly on the top outlier patch. These maps near-perfectly align with the high-norm outliers ("Patch Norms").
@@ 3.2 Register neurons
1:Input: Image set $\mathcal{I}=\{I_{1},\dots,I_{M}\}$, maximum layer index top_layertop\_layer, number of register neurons to return top_ktop\_k, number of neurons in each layer NN
2:Output: top_ktop\_k register neurons
@@
4:for all $I_{i}\in\mathcal{I}$ do
5:  $O\leftarrow\mathrm{FindOutliers}(I_{i})$ # get indices of top-norm patches
6:  for $\ell=0$ to top_layertop\_layer do
7:    for $n=0$ to N1N-1 do
8:     $\textit{avg\_act}[\ell,n]\leftarrow\textit{avg\_act}[\ell,n]+\frac{1}{|O|M}\sum_{p\in O}\mathrm{activation}_{\ell,n}(I_{i},p)$
@@
10:  end for
11:end for
12:return top_ktop\_k neurons with largest avg_act
Detecting register neurons. Based on our hypothesis, we formulate a simple algorithm to find register neurons in Algorithm˜1. Our method finds neurons whose average activation at outlier positions is consistently high across images. To detect outlier positions (FindOutliers within Algorithm˜1), we follow Darcet et al. (2024) and output the positions for which the norms of the corresponding image tokens are above a predefined threshold. Our algorithm searches for neurons in preceding layers before outliers start, set by the top_layertop\_layer parameter. We output top_ktop\_k neurons with the highest average activations. We present a discussion of hyperparameters in Section˜A.4. For OpenCLIP, we set top_layer=5top\_layer=5, the outlier threshold at 75, and top_k=10top\_k=10. Since register neurons determine where outliers emerge, we can also intervene upon them to move outliers to arbitrary positions, as demonstrated next.
Moving outliers with register neurons. To demonstrate the importance of these register neurons, we can use them to “move” outliers to different image locations. We first apply FindRegisterNeurons to detect the register neurons. We then modify their activation pattern during the forward pass – for each register neuron, we copy the highest neuron activation across the tokens into the locations of the tokens to which we want to move the outliers. We zero out the activations of the neuron elsewhere.
Register neurons causally set the position of outliers. To test our ability to move outliers to arbitrary spots, we assign the highest activation to a random patch and measure its last-layer output norm, the maximum norm of other patches, and the highest last-layer [CLS] attention, both to the selected outlier and any other patch. Successful interventions yield high norms and attention for the targeted patch, with low values elsewhere. As shown in Figure˜4, modifying the activations of register neurons effectively controls where outliers emerge. In contrast, intervening on random neurons has little effect (Section˜A.11.3). Figure˜1 demonstrates that intervening on register neurons can make outliers appear in various counts and spatial patterns (e.g. a heart). We use this technique to mimic registers, as follows in the next section.
Figure 4: Intervening on activations of register neurons effectively shifts outliers to random patches and test-time registers. For all register neurons, we copy their highest activation into a selected patch and zero out the activations elsewhere. Left: norm of chosen random patch (yellow) and max norm of any other patch (blue). Right: [CLS] attention to chosen random patch (yellow) and max [CLS] attention (blue) to any other patch. Our intervention can shift outliers to randomly selected patches as well as test-time registers (see Section˜A.11.2).
@@ 4 Adding registers at test-time
Moving outliers to added tokens. As outlier patches lose their local patch information (Darcet et al., 2024), it is undesirable to have them within the image since it may affect downstream performance. Previous work has suggested retraining ViTs with extra tokens to remove high-norm artifacts and attention sinks. However, retraining existing ViTs is expensive, so we propose to add an extra input token and move the outliers there with register neurons. Algorithm 2 (Section˜A.3), with accompanying visualization in Figure˜11, specifies our test-time register edit. At each register neuron, we copy the maximum patch activation to the register token, set all other token activations to zero, and resume the forward computation. Importantly, these outliers are essential to the model’s internal computations and must appear within a token. For instance, zeroing out the register neurons’ activation maps instead of moving the outliers causes OpenCLIP ViT-B/16’s zero-shot ImageNet performance to drop $\sim$15%. Ablating a set of random neurons results in minimal drop: $69.3\%\pm 1.1.$
Test-time register initialization. We initialize our extra token to be a vector of zeros. We assess several initialization strategies, but we find that they do not significantly affect the ability of test-time registers to store the high norms (Section˜A.6). Additionally, we focus our investigation on using one test-time register and report the impact of using more registers in Section˜A.5.
Outliers move outside the image after adding test-time registers. To evaluate whether test-time registers can absorb the outliers, we apply our intervention and measure the max norms of image patch outputs and the test-time register. Our intervention results are nearly identical to the distribution of patch norms and [CLS] attention after shifting outliers to random patches, previously shown in Figure˜4 (see results for test-time registers in Section˜A.11.2). The image patches no longer have outliers, whereas the test-time register absorbs the outlier norms. This change is also present in the attention maps (Figure˜5), which no longer have noisy artifacts and match the quality of attention maps from models with trained registers. [^2] As there is no open-source OpenCLIP trained with registers, we only present comparisons on DINOv2. See Section˜A.9 for attention maps in all model layers.
Test-time registers hold global information. We verify that test-time registers encapsulate global image information (e.g., image class) similarly to trained registers (Darcet et al., 2024). To assess this, we perform linear probing on both trained and test-time registers for classification on ImageNet (Deng et al., 2009), CIFAR-10, and CIFAR-100 (Krizhevsky et al., 2009). We also compare their performance to the [CLS] token and a token that corresponds to a central patch in the image. As shown in Table˜1, the classification accuracy of test-time registers closely matches that of trained registers and is slightly lower than the [CLS] token performance. These results suggest that test-time registers, like their learned counterparts, effectively capture global image-level information.
Figure 5: Qualitative results on attention maps w/ test-time registers. We present the last layer’s mean [CLS] attention maps in DINOv2 and compare them to the model with trained registers. Test-time registers produce similarly high-quality maps as trained registers.